a) Complete the statement of: Stokes Theorem: Let S be an oriented surface bounded by a piecewise smooth simple closed curveb) Use Stokes theorem to write as an iterated integral, J. (y, -2, 1)odr where is the circle of radius 1 in the plane z = 1
Answer #1

@ stokex theorem. then, 07.027 = 1/(VX) - ads 5 Where nis the outward drown unit vector. direction of cis to taken for whichNow we write, II (V x F ) . ñ ds 2xê tagit i i Allere sigium as z=2-2²-4², where xty²21 n = v( 2-2 + x² + y2), = ax i tay p t